OpenStudy (anonymous):

find the product write in simplest form 6/12x7/9

OpenStudy (kc_kennylau):

do you have any idea how to simplify \(\dfrac6{12}\)?

OpenStudy (anonymous):

its hard for me:/

OpenStudy (kc_kennylau):

has it occurred to you that 12 is 2 times 6?

OpenStudy (anonymous):

I know that

OpenStudy (kc_kennylau):

and that \(\dfrac6{12}\) is equivalent to \(\dfrac12\)?

OpenStudy (kc_kennylau):

So we have: \[\frac6{12}\times\frac79\] Cancelling 6 we have: \[\frac12\times\frac79\] Multiplying the denominators we have: \[\frac7{18}\]

OpenStudy (anonymous):

6/12 × 7/9 = (6 × 7) / (12 × 9) = 42/108 = 7/18